OpenStudy (haleyelizabeth2017):

Write the polar equation in rectangular form. r cos theta=5

rebeccaxhawaii (rebeccaxhawaii):

z=a+jb

OpenStudy (haleyelizabeth2017):

eh?

OpenStudy (dumbcow):

\[x^2 + y^2 = r^2\] \[\tan \theta = \frac{y}{x}\] Note also identity: \[\sec^2 \theta = 1 + \tan^2 \theta\]

OpenStudy (dumbcow):

rewrite equation \[r = 5 \sec \theta\] square both sides \[r^2 = 25 \sec^2 \theta\] substitute \[x^2 +y^2 = 25(1 + \frac{y^2}{x^2})\]

OpenStudy (haleyelizabeth2017):

You lost me...

OpenStudy (dumbcow):

which part

OpenStudy (haleyelizabeth2017):

How did you end up with tan?

OpenStudy (haleyelizabeth2017):

In order to get r isolated, you divided both sides by cos theta, correct?

OpenStudy (dumbcow):

from the identity sec^2 = 1 +tan^2

OpenStudy (dumbcow):

yes

OpenStudy (haleyelizabeth2017):

that would have given us \[\frac{5}{cos \theta}\]which is also...what?

OpenStudy (haleyelizabeth2017):

Wait...it's also sec -_-

OpenStudy (haleyelizabeth2017):

so that's how you got 5sec... I gotchya now ;)

OpenStudy (dumbcow):

wait hold on ... i forgot something i made this harder than it really is \[x = rcos \theta\] \[y = r \sin \theta\] all you have to do is sub in "x" for rcos .... leaving \[x = 5\]

OpenStudy (haleyelizabeth2017):

lol I got what you meant in the end after I went through it XD

OpenStudy (dumbcow):

ok good :)

OpenStudy (haleyelizabeth2017):

Thank you very much (: